Find the sum of vector A and C
A=8i + 8j
C=8i - 16j
1
Expert's answer
2020-02-26T10:12:14-0500
"\bold{A+C}=(8\bold{i}+8\bold{j})+(8\bold{i}-16\bold{j})"


"\bold{A+C}=(8+8)\bold{i}+(8-16)\bold{j}"


"\bold{A+C}=16\bold{i}-8\bold{j}"